# Non-Heisenbergian Quantum Mechanics

### v2

@@ V Interacting particles system
Finally, it is important to remark that the corrected energy levels in our framework differ from the ones in the deformed commutation relation models Stetsko and Tkachuk (2006); Bouaziz and Ferkous (2010); Quesne and Tkachuk (2010). Especially, in contrast to our result, in some deformed commutation relation models, the hydrogen energy spectrum depends on the center-of-mass momentum Quesne and Tkachuk (2010). In principle, these differences can be used to experimentally distinguish between these models and our non-Heisenbergian quantum mechanics.
VI Summary and out-looks
@@
Understanding the core content of quantum mechanics requires us to disentangle the hidden logical relationships between the basics assumptions and results of this theory. One of the most important assumptions of Heisenberg quantum mechanics is that there exists a position self-adjoint operator that satisfies the canonical commutation relation. This assumption is usually considered necessary to derive most of the important results of the theory, especially the uncertainty relation. Nonetheless, in this work, by ignoring the position operator, we demonstrate that Heisenberg quantum mechanics could be seen as a special case of a more general self-consistent quantum theory, which has most of fundamental characteristics of the Heisenberg quantum mechanics, such as the Galilean symmetry, probability conservation, position-momentum uncertainty relation, the discrete and bounded from bellow energy spectrum in bound states, and so on. This approach opens an important door to generalize the quantum theory consistent with a given space-time symmetry. It is important to remark that the deformation of commutation relation, as used in GUP literature Maggiore (1993b); Kempf et al. (1995); Fadel and Maggiore (2022), is not consistent with usual space-time symmetries Fadel and Maggiore (2022); Maggiore (1994); Bosso et al. (2023a)—the usual canonical commutation relation is a direct result of the Galilean and also Poincare algebras if we consider a self adjoint position operator that correctly transform under special translation. Therefore, it seems that for constructing a non-relativistic quantum theory with a non-zero minimal position uncertainty, the presented formalism is more acceptable than deformed commutation relation approach.
@@
One of the most important extensions of the present work is its relativistic extension. It could provide a way to circumvent the fundamental problems in relativistic quantum mechanics, which arise from the impossibility of defining a completely proper relativistic position operator. In fact, due to the problems mentioned in the Introduction and some other issues Sebens (2019); Kazemi et al. (2018); Terno (2014), despite much effort in recent years Kowalski and Rembieliński (2011); Wagner et al. (2011, 2012); Bialynicki-Birula and Bialynicka-Birula (2012); Babaei and Mostafazadeh (2017); Kazemi et al. (2018); Kiessling and Tahvildar-Zadeh (2018); Sebens (2019); Hawton (2019, 2021); Glasgow and Ware (2020), a complete and agreed-upon formulation of relativistic quantum mechanics with a clear space-time resolution has not yet been proposed. Nonetheless, in relativistic quantum experiments, the position measurements are routinely performed with high accuracy (just for example see Sinha et al. (2010); Pleinert et al. (2021); Korzh et al. (2020)), and thus, a complete quantum theory must be able to predict the observed position distributions Sebens (2019); Krekora et al. (2004). We believe that a proper relativistic extension of our formulation of quantum theory without a position operator could circumvent the mentioned problems, as has been previously shown by one of the authors in a very special case Kazemi et al. (2018). The detailed study of the relativistic extension is left for future works.
@@
Finally, it should be noted that the generalization of quantum theory without a position operator is natural and interesting from the perspective of some non-standard interpretations of quantum theory, such as Bohmian mechanics Barnea et al. (2018), Nelson stochastic mechanics Nelson (1966), and the many interacting worlds model Hall et al. (2014). In these theories, the operators are not fundamental; they are secondarily mathematical objects that are useful for computing statistical observable results Dürr et al. (2004a). However, in the formalism of these theories, the position probability density and probability current have fundamental important roles. So, the approach of the present work could shed some new light on the generalization of these alternative formulations of quantum theory, especially in the relativistic regime, which is an interesting open problem Foo et al. (2022); Dürr et al. (2014); Gisin (2018); Holland (1995); Dürr et al. (2004b).
VII Acknowledgment

### v4

@@ 5 Interacting particles system
Finally, it is important to remark that the corrected energy levels in our framework differ from the ones in the deformed commutation relation models [88, 89, 90, 91]. In the first order of perturbation, the conventional GUP model, based on deformed commutation relations, yields the following energy spectrum correction for the hydrogen atom [91]:
$\Delta E_{n}=\frac{e^{2}}{2a_{0}n^{2}}$ $\left[\frac{2}{3}\frac{l_{1}^{2}m_{1}+l_{2}^{2}m_{2}}{a_{0}^{2}(m_{1}+m_{2})}\frac{8n-3}{n^{2}}\right$.
$+\left.\frac{20}{9\hbar^{2}}\frac{l_{1}^{2}m_{1}+l_{2}^{2}m_{2}}{m_{1}+m_{2}}m_{1}m_{2}v^{2}\right],$
where $v$ denotes the center-of-mass velocity of the hydrogen atom. In particular, unlike the result of our model, the energy spectrum of the hydrogen atom depends on its center-of-mass velocity. In principle, this difference, which arises as a consequence of Galilean symmetry breaking in the conventional GUP model, can be used to distinguish experimentally between this model and our model.
6 Summary and out-looks
@@
A clear understanding of the foundational structure of quantum mechanics requires disentangling the implicit logical connections between its basic assumptions and derived results. One of the most important assumptions of Heisenberg’s quantum mechanics is the existence of a self-adjoint position operator that satisfies the canonical commutation relation. This assumption is usually considered necessary to derive most of the important results of the theory, especially the uncertainty relation. Nonetheless, in this work, by ignoring the position operator, we demonstrate that Heisenberg quantum mechanics can be interpreted as a special case of a more general self-consistent quantum theory—one that retains most of the fundamental characteristics of Heisenberg’s formulation, including Galilean symmetry, probability conservation, the position-momentum uncertainty relation, a discrete and bounded-from-below energy spectrum for bound states, and so on.
In this work, the POVM formalism is used to describe the intrinsic minimal spatial resolution at the fundamental level, not to describe the effective resolution of the measurement apparatus, which has usually been studied in the literature. This is not merely a conceptual reinterpretation of existing formalism. Our interpretation leads to new dynamics and so distinct phenomenological predictions. Note that when POVMs are employed merely to model limitations of apparatus resolution, the underlying Hamiltonian remains unaffected. In contrast, assuming an intrinsic minimal spatial resolution implies that the position operator is no longer well-defined at the fundamental level. As a result, the Hamiltonian must be constructed directly from the position POVM, rather than from the position operator. This leads to a modification of the dynamics—a feature absent in previous works and one that emerges naturally from our distinct physical interpretation.
It is important to stress that our modified dynamic is fundamentally different from those of conventional GUP frameworks that rely on deformed commutation relations [55, 56, 57]. This difference leads to distinct physical consequences at the phenomenological level. From a foundational perspective, a key distinction lies in the fact that our model is developed in full accordance with the standard non-relativistic space-time symmetries. In contrast, the framework of deformed commutation relations tends to break such symmetries [57, 58, 59]—The usual canonical commutation relation is a direct result of the Galilean and also Poincare algebras if we consider a self-adjoint position operator that correctly transforms under spatial translation. Therefore, it appears that our formalism is more suitable than the deformed commutation relation approach for constructing a non-relativistic quantum theory with a non-zero minimal position uncertainty.
It is valuable to explore a natural extension of the present study. A discerning reader may question whether both the canonical position and momentum operators could be substituted with POVMs. Would such a substitution yield a quantum theory aligned with the extended generalized uncertainty principle (EGUP) [92, 93, 94, 61]? In Appendix B, we address these questions in detail and demonstrate that, at least within the framework of Galilean symmetry, this approach does not yield a theory more general than the model proposed in this paper.
@@
One of the most important extensions of the present work is its relativistic extension. This could provide a way to circumvent the fundamental problems in relativistic quantum mechanics, which arise from the impossibility of defining a proper relativistic position operator. In fact, due to the problems mentioned in the Introduction and some other issues [95, 16, 6], despite much effort in recent years [96, 29, 30, 32, 97, 16, 98, 95, 99, 100, 31], a complete and agreed-upon formulation of relativistic quantum mechanics with a clear space-time resolution has not yet been proposed. Nonetheless, in relativistic quantum experiments, the position measurements are routinely performed with high accuracy (for example see [25, 26, 27]), and thus, a complete quantum theory must be able to predict the observed position distributions [95, 15]. We believe that a proper relativistic extension of our formulation of quantum theory without a position operator could circumvent the aforementioned problems, as has been previously shown by one of the authors in a very special case [16]. The detailed study of the relativistic extension is left for future works.
@@
It should be noted that the generalization of quantum theory without a position operator is natural and interesting from the perspective of some non-standard interpretations of quantum theory, such as Bohmian mechanics [101], Nelson stochastic mechanics [102], and the many interacting worlds model [103]. In these theories, the operators are not fundamental; rather, they are secondary mathematical objects useful for computing statistically observable results [104]. However, in the formalism of these theories, the position probability density and probability current have fundamental roles. Therefore, the approach of the present work could shed new light on the generalization of these alternative formulations of quantum theory, especially in the relativistic regime, which is an interesting open problem [105, 106, 107, 108, 109].
7 Acknowledgment